# Self-Supervised Multisensory Pretraining for Contact-Rich Robot Reinforcement Learning

### v1

@@ V-A Baselines and Training Details
We compare our methods against one transformer and two non transformer-based baselines. The latter models extract sensor-specific features with a CNN for vision and an MLP for proprioception and force torque readings. The Concat model fuses the concatenated features with a 2-layer MLP to form the multisensory latent representation [lee_making_2018]. The PoE model generates separate means and variances for each sensor and fuses them with a product of experts approach. VTT [chen_visuo-tactile_2022] fuses all sensors via a transformer-encoder and compresses the features via multiple linear layers to a compact latent. As the baseline models introduce a bottleneck, we neglect masking and pretrain the encoders with observation reconstruction to form the latent. All models use a similar decoder architecture for the representation objectives, where proprioception and force torque signals are reconstructed using separate two-layer MLPs and vision via a deconvolutional CNN.
@@ V-B Multisensory Environments
Our multisensory environments are based on panda-gym [gallouedec_panda-gym_2021] and Pybullet [coumans_pybullet_2016]. The tasks Peg Insertion, Push Cube, Close Drawer Gently and Dual Arm Peg Insertion are challenging contact-rich manipulation tasks with sensor noise, varying dynamics and number of sensors. Each task has a dense reward function and comes with multisensory observations consisting of proprioception (joint position and velocities), vision from an external RGB camera and a force torque sensor located on the wrist of the robot. The observation of the Dual Arm environment has an additional FT-sensor and proprioception of the second robot arm. The action space consists of the 3-dimensional end-effector displacement. Task details are provided below and an overview of the environments can be found in Figure 3.
@@ V-C Simulation Experiments
Figure 4 shows the performance of our MSDP-P and MSDP-R models compared to the baselines. We obtain superior performance in Peg Insertion and Push Cube, while also outperforming in Close Drawer Gently and Dual Arm Peg Insertion, indicating a expressive representation for contact-rich tasks. Especially in Peg Insertion, our sensor fusion provides fine-grained features to solve the task around 80% of the time after only $200,000$ environment steps. The baselines struggle to combine sensor knowledge to consistently insert the peg. Baseline models need an extended amount of training in Push Cube to be able to account for varying object dynamics, like mass and center of mass of the cube, while our representations allow quick adaptation to those changes, leading to a steep learning curve and optimal performance. The Close Drawer Gently task offers no clear results, mainly due to simulation bottlenecks. We observed sudden changes in the drawer during contact that limit the agent’s ability to close it smoothly. Still, our method can learn the task faster with better final performance. For the Dual Arm Peg Insertion task additional sensor encoders are used to map the proprioceptive state of the second robot and its force torque reading to the respective latent. Next to PoE we achieve optimal performance over all runs. Our embedding-based latent would allow for more advanced mapping to actor/critic, which we omit for a fair comparison.
Sensor Ablation: To evaluate the role of individual sensors, we conducted sensor ablation studies on Peg Insertion and Push Cube using the MSDP-P model. Figure 5 compares the success rate in various sensor settings, highlighting the performance gains compared to a vision only approach. Proprioception is essential to identify the peg pose under vision noise, while the force torque sensor guides exploration and allows for consistent insertions, indicated by the lower confidence interval. Figure 6 shows the episode lengths in the Push Cube task as all sensor combinations achieve near-optimal performance. Vision is sufficient to locate the cube and solve the task. The usage of all sensor allows the agent to make and keep contact with the cube to solve the task in an efficient manner, indicated by the fast completion time (30% faster compared to vision only). In addition to the performance gains, we observe smoother object interactions with access to the force torque modality. Most prominent in the Peg Insertion task, where the agent without having access to the FT-sensor asserts high forces on the hole, indicating unwanted exploration actions which may damage the robot or environment in the real world.
@@
Latent Bridging: The latent bridging mechanism to obtain a compact representation from multisensory embeddings, described in Section IV-C, can have a substantial impact on performance. We compare our cross-attention extraction with common approaches. CLS is using the CLS embedding of the encoder, commonly used in the Vision Transformer [dosovitskiy_image_2020] or Imitation Learning [saxena_mrest_2024]. In Pooling we mean-pool all sensor embeddings similar to [sferrazza_power_2023, seo_masked_2023]. To account for the different number of embeddings between sensors, we first take the mean of all vision embeddings. To avoid possible dilution Cat provides the Concatenation of sensor embeddings, where we again pool the vision embeddings. Common latent bridging mechanisms do not introduce new learnable parameters, while our cross-attention only adds minimal parameters to extract task-specific features.
Figure 7: Push Cube latent bridging ablation: Our mechanism extracts fine grained multisensory features resulting in efficient policy training.
@@ V-D Real World Experiments
We conduct the Peg Insertion and Push Cube tasks in the real world using a Franka robot arm with a wrist-mounted FT-sensor and a custom end-effector. The observation space consists of endeffector position and velocity as proprioception, four force torque readings and an downsampled 64 by 64 RGB image from the 3rd person RealSense camera for both tasks. Peg and hole are both 3d printed. The robot needs to align the orientation of the triangular peg and place it precisely to fully insert the peg and complete the task. In Push Cube the agent needs to push the 7.5 cm block to the 15 cm away goal location. Successful episodes are detected with the endeffector position in Peg Insertion and an aruco marker and a second camera in the Push Cube task. The starting position of the robot is randomized and we use a sparse reward of +1 upon success in both tasks. An overview of the setup and results can be found in Figure 11(a) and 11(b).
We build upon the SERL-package [luo2024serl] and use a cartesian impedance controller for safe interaction. The multisensory encoder is pretrained with data from 20 demonstrations and ~2,000 samples of play data, resulting in a total of ~3,000 samples. Downstream tasks are trained directly on the real system for only 6,000 environment interactions with the proposed RLPD [ball_efficient_2023] algorithm from SERL. We pretrain for 6,000 steps in Peg Insertion, updating critic and actor twice per online sample. For Push Cube, due to higher vision variance, we use 10,000 update steps and apply four critic and one actor update step per online interaction to obtain meaningful representations and efficient training. The complete training pipeline, including data collection, pretraining, and online RL, learns the final policy in less than 55 minutes for both tasks.
@@
Using online available datasets for pretraining has been omitted in this work, due to limited multisensory datasets and a large domain gap. We will open-source all simulation environments and our multisensory robotic dataset upon publication to facilitate future work.
@@ VI Conclusion and Future Work
In this work, we propose MultiSensory Dynamic Pretraining (MSDP), a novel pretraining framework to learn multisensory representations for contact-rich manipulation tasks. Specifically, MSDP learns to reconstruct varying sensory information from a subset of input sensor embeddings, leading to cross-sensor predictions and sensor fusion. MSDP captures the interplay between different sensor modalities to learn a rich multisensory representation, which in combination with task-specific feature extraction through cross-attention, leads to a superior performance in challenging contact-rich manipulation tasks in simulation and the real world. Future work may investigate MSDP’s capabilities for sensor modalities beyond vision and force-torque, such as tactile or audio. Finally, we reckon that extending MSDP’s pretraining stage to incorporate play data from environments at large scale may yield representations that generalize over new tasks.
VII Acknowledgement
This research is funded by the German Research Foundation (DFG) Emmy Noether Programme (CH 2676/1-1), the EU’s Horizon Europe project “ARISE” (Grant no.: 101135959) and the German Federal Ministry of Education and Research (BMBF) project Robotics Institute Germany (RiG) (Grant no.: 16ME1001).

### v2

@@ V-A Baselines and Training Details
We compare our methods against one transformer and two non transformer-based baselines. The latter models extract sensor-specific features with a CNN for vision and an MLP for proprioception and force torque readings. The Concat model fuses the concatenated features with a 2-layer MLP to form the multisensory latent representation [26]. The PoE model generates separate means and variances for each sensor and fuses them with a product of experts approach. VTT [6] fuses all sensors via a transformer-encoder and compresses the features via multiple linear layers to a compact latent. As the baseline models introduce a bottleneck, we neglect masking and pretrain the encoders with observation reconstruction to form the latent. All models use a similar decoder architecture for the representation objectives, where proprioception and force torque signals are reconstructed using separate two-layer MLPs and vision via a deconvolutional CNN.
@@ V-B Multisensory Environments
Our multisensory environments are based on panda-gym [12] and Pybullet [7]. The tasks Peg Insertion, Push Cube, Close Drawer Gently and Dual Arm Peg Insertion are challenging contact-rich manipulation tasks with sensor noise, varying dynamics and number of sensors. Each task has a dense reward function and comes with multisensory observations consisting of proprioception, vision and force torque readings. The observation of the Dual Arm environment has an additional FT-sensor and proprioception of the second robot arm. The action space consists of the 3-dimensional endeffector displacement. Task details are provided below and an overview of the environments can be found in Figure 3.
@@ V-C Simulation Experiments
Figure 4 shows the performance of our MSDP-P and MSDP-R models compared to the baselines. We obtain superior performance in Peg Insertion and Push Cube, while also outperforming in Close Drawer Gently and Dual Arm Peg Insertion, indicating an expressive representation for contact-rich tasks. Especially in Peg Insertion, our sensor fusion provides fine-grained features to solve the task around 80 % of the time after only $200,000$ environment steps. The baselines struggle to combine sensor knowledge to consistently insert the peg. Baseline models need an extended amount of training in Push Cube to be able to account for varying object dynamics, like mass and center of mass of the cube, while our representations allow quick adaptation to those changes, leading to a steep learning curve and optimal performance. The Close Drawer Gently task offers no clear results, mainly due to simulation bottlenecks. We observed sudden changes in the drawer during contact that limit the agent’s ability to close it smoothly. Still, our method can learn the task faster with better final performance. For the Dual Arm Peg Insertion task additional sensor encoders are used to map the proprioceptive state of the second robot and its force torque reading to the respective latent. Next to PoE we achieve optimal performance over all runs. Our embedding-based latent would allow for more advanced mapping to actor/critic, which we omit for a fair comparison.
@@
Sensor Ablation: To evaluate the role of individual sensors, we conduct sensor ablation studies on Peg Insertion and Push Cube using the MSDP-P model. Figure 5 compares the success rate in various sensor settings, highlighting the performance gains compared to a vision only approach. Proprioception is essential to identify the peg pose under vision noise, while the force torque sensor guides exploration and allows for consistent insertions, indicated by the lower confidence interval. Figure 6 shows the episode lengths in the Push Cube task as all sensor combinations achieve near-optimal performance. Vision is sufficient to locate the cube and solve the task. The usage of all sensor allows the agent to make and keep contact with the cube to solve the task in an efficient manner, indicated by the fast completion time (30 % faster compared to vision only). In addition to the performance gains, we observe smoother object interactions with access to the force torque modality. Most prominent in the Peg Insertion task, where the agent without having access to the FT-sensor asserts high forces on the hole, indicating unwanted exploration actions which may damage the robot or environment in the real world.
@@
Latent Bridging: The latent bridging mechanism to obtain a compact representation from multisensory embeddings, described in Section IV-C, can have a substantial impact on performance. We compare our cross-attention extraction with common approaches. CLS is using the CLS embedding of the encoder, commonly used in the Vision Transformer [9] or Imitation Learning [42]. In Pooling we mean-pool all sensor embeddings similar to [44, 43]. To account for the different number of embeddings between sensors, we first take the mean of all vision embeddings. To avoid possible dilution Cat provides the Concatenation of sensor embeddings, where we again pool the vision embeddings. Common latent bridging mechanisms do not introduce new learnable parameters, while our cross-attention only adds minimal parameters to extract task-specific features.
Figure 7: Push Cube latent bridging ablation: Our mechanism extracts fine-grained multisensory features resulting in efficient policy training.
@@ V-D Real World Experiments
We conduct the Peg Insertion and Push Cube tasks in the real world using a Franka robot arm with a wrist-mounted FT-sensor and a custom endeffector. The observation space consists of endeffector position and velocity as proprioception, four force torque readings and an downsampled 64 by 64 RGB image from the 3rd person RealSense camera for both tasks. Peg and hole are both 3d printed. The robot needs to align the orientation of the triangular peg and place it precisely to fully insert the peg and complete the task. In Push Cube the agent needs to push the 7.5 cm block to the 15 cm away goal location. Successful episodes are detected with the endeffector position in Peg Insertion and an aruco marker and a second camera in the Push Cube task. The starting position of the robot is randomized and we use a sparse reward of +1 upon success in both tasks. An overview of the setup and results can be found in Figure 11(a) and 11(b).
We build upon the SERL-package [34] and use a cartesian impedance controller for safe interaction. The multisensory encoder is pretrained with data from 20 demonstrations and ~2,000 samples of play data, resulting in a total of ~3,000 samples. Downstream tasks are trained directly on the real system for only 6,000 environment interactions with the proposed RLPD [3] algorithm from SERL. We pretrain for 6,000 steps in Peg Insertion, updating critic and actor twice per online sample. For Push Cube, due to higher vision variance, we use 10,000 update steps and apply four critic and one actor update step per online interaction to obtain meaningful representations and efficient training. The complete training pipeline, including data collection, pretraining, and online RL, learns the final policy in less than 55 minutes for both tasks. Compared to prior real-world multisensory RL methods [26, 6], our approach requires only a fraction of the data (3 % pretraining data in Peg Insertion) and training time (< 20 %) to learn the final policy.
@@
Robustness Evaluation: We evaluate the final policy of MSDP-P in the Peg Insertion task under various disturbances to showcase the robustness of our pretrained multisensory encoder and policy. We evaluate each change that hasn’t been observed during training for 20 trials. Trained on $K_{c}=2000$ cartesian-stiffness the policy achieves 90 % success rate with decreased ($K_{c}=1500$) and 100 % with increased cartesian-stiffness ($K_{c}=2500$). MSDP shows remarkable robustness against changed light settings, e.g., back light (100 %), front light (100 %), disco lights (100 %) and visual occlusion (partly blocked camera view, 95 %) and external forces (80 %). All visual disturbances are shown in Figure 12.
Figure 12: Visual disturbances of the robustness evaluation. From left to right: back light, front light, occlusion, and disco lights. MSDP achieves a high success rate across all disturbances, which have not been seen during training.
This work assumes the presence of all sensor modalities at the start of training to obtain an expressive representation from limited multisensory data. The usage of MSDP to enhance existing representations seems promising to leverage common available datasets for generalization.
@@ VI Conclusion and Future Work
In this work, we propose MultiSensory Dynamic Pretraining (MSDP), a novel pretraining framework to learn multisensory representations for contact-rich manipulation tasks. Specifically, MSDP learns to reconstruct varying sensory information from a subset of input sensor embeddings, leading to cross-sensor predictions and sensor fusion. MSDP captures the interplay between different sensor modalities to learn a rich multisensory representation, which in combination with task-specific feature extraction through cross-attention, leads to a superior performance in challenging contact-rich manipulation tasks in simulation and the real world. Future work may investigate MSDP’s capabilities for sensor modalities beyond vision and force-torque, such as tactile or audio, as well as its use for fine-tuning existing representations with multisensory data. Finally, we reckon that extending MSDP’s pretraining stage to incorporate play data from environments at large scale may yield representations that generalize over new tasks.